### v2

# Revisiting the Thickness of the Air-Water Interface from Two Extremes of Interface Hydrogen Bond Dynamics

Abstract
The air-water interface plays a crucial role in many aspects of science, because of its unique properties, such as a two-dimensional hydrogen bond (HB) network and completely different HB dynamics compared to bulk water. However, accurately determining the boundary of interfacial and bulk water, that is, the thickness of the air-water interface, still challenges experimentalists. Various simulation-based methods have been developed to estimate the thickness, converging on a range of approximately 3 to 10 Å. In this study, we introduce a novel approach, grounded in density functional theory-based molecular dynamics (DFTMD) simulations, to measure the air-water interface thickness, offering a different perspective based on prior research. To capture realistic HB dynamics in the air-water interface, two extreme scenarios of the interface HB dynamics are obtained: one underestimates the interface HB dynamics, while the other overestimates it. Surprisingly, our results suggest that the interface HB dynamics in both scenarios converge as the thickness of the air-water interface increases to 4.0 Å. This convergence point, indicative of the realistic interface thickness, is confirmed by our calculation of anisotropic decay of OH stretch at the air-water interface and supported by the results from recent ab initio molecular dynamics simulations.
preprint: AIP/123-QED
I Introduction
The air-water interface has been the subject of extensive study due to its ubiquity in nature and its unusual macroscopic properties as a model system for aqueous hydrophobic interfaces Morita and Hynes (2000); Jungwirth and Tobias (2001); Wilson et al. (2002); Raymond et al. (2003); Shen and Ostroverkhov (2006); Smits et al. (2007); Sovago et al. (2008, 2009); Stiopkin et al. (2011); Hsieh et al. (2011); Nihonyanagi et al. (2011); Singh et al. (2014); Nihonyanagi et al. (2015); Ishiyama (2021); Ahmed, Nihonyanagi, and Tahara (2022); Omranpour et al. (2024). It is widely accepted that water molecules behave in a completely different manner at the interface than in the bulk phase Lee, McCammon, and Rossky (1984); Dang and Chang (1997a); Hsieh et al. (2013).
Advances in the study of hydrogen bond (HB) dynamics at the air-water interface have been significant. Liu et al. Liu, Harder, and Berne (2005) used molecular dynamics (MD) simulations to demonstrate faster HB breaking and forming at the interface than bulk water, attributed to quicker translational diffusion. From sum frequency generation (SFG) vibrational spectroscopy, Gan et al. Gan et al. (2006a) found that at the air-water interface, singly hydrogen (H)-bonded water molecules align almost parallel to the interface with limited orientational variation, while doubly H-bonded donor molecules orient their dipole vectors away from the liquid phase, highlighting diverse behaviors among interfacial water molecules. Almost concurrently, through time-resolved SFG vibrational spectroscopy, McGuire and ShenMcguire and Shen (2006) observed ultrafast vibrational dynamics at the interface, noting that the relaxation behaviors of interfacially bonded OH stretch modes on sub-picosecond timescales were akin to those in bulk water, encompassing spectral diffusion, vibrational relaxation, and thermalization. Pioneering work by Tahara’s group Singh et al. (2013); Inoue et al. (2016), which presented the first two-dimensional heterodyne-detected vibrational SFG (2D HD-VSFG) spectra of the OH stretch region at the interface, highlighted diverse behaviors among interfacial HB OH groups. Subsequent studies, including those by Jeon et al. Jeon et al. (2017) and Ojha and Kühne Ojha and Kühne (2021), have used MD and ab initio MD (AIMD) simulations to explore the structure and dynamics of interfacial water, uncovering weaker H-bonds and faster vibrational spectral dynamics of free OH groups compared to H-bonded OH groups at the interface. These collective insights enhance our understanding of the vibrational energy relaxation, HB dynamics, and interactions of water molecules at the air-water interface. Building upon this knowledge of interfacial behavior, significant efforts have also been directed towards quantifying the physical characteristics of the interface, including its thickness, which plays a crucial role in understanding its molecular interactions and behavior.
The air-water interface thickness has been measured via ellipsometry Raman and Ramdas (1927); McBain, Bacon, and Bruce (1939); Kinosita and Yokota (1965), relative permittivity measurements Fumagalli et al. (2018), X-ray reflectivity Braslau et al. (1985, 1988), SFG spectroscopy Superfine, Huang, and Shen (1991); Du et al. (1993); Ostroverkhov, Waychunas, and Shen (2005); Groenzin et al. (2007); Shen (2012); Ishiyama, Takahashi, and Morita (2012); Shen (2013); Singh et al. (2013); Ohto et al. (2015); Tang et al. (2020), classical MD simulationsLiu, Harder, and Berne (2005); Kuo et al. (2006); Wick et al. (2007); Morita and Garrett (2008); Sedlmeier, Horinek, and Netz (2009); Fan et al. (2009); Ishiyama and Morita (2009); Willard and Chandler (2010); Vila Verde, Bolhuis, and Campen (2012); Ishiyama, Morita, and Tahara (2015); Nagata et al. (2015) and the AIMD simulationsKühne et al. (2011); Nagata et al. (2012); Kessler et al. (2015); Jeon et al. (2017); Ohto et al. (2019); Tang et al. (2020) to mention just a few. There is a consensus that the thickness of the air-water interface is about 3 to 10 Å Jungwirth and Tobias (2006); Sedlmeier, Horinek, and Netz (2009); Ishiyama and Morita (2009); Kuo et al. (2006); Wick et al. (2007); Fan et al. (2009); Kühne et al. (2011); Baer et al. (2014); Kessler et al. (2015); Nguyen, Nakahara, and Phan (2020); Peng et al. (2021). Nonetheless, accurately determining the thickness remains experimentally challenging. The MD and Monte Carlo (MC) simulations of the air-water interface yield molecular-level information not readily available in experiments. These simulations, which utilize various intermolecular potential functions, have played a crucial role in estimating the thickness Townsend, Gryko, and Rice (1985); Wilson, Pohorille, and Pratt (1987); Matsumoto and Kataoka (1988); Townsend and Rice (1991); Lie et al. (1993); Alejandre, Tildesley, and Chapela (1995); Taylor, Dang, and Garrett (1996); Dang and Chang (1997b); Tang et al. (2020). Additionally, density functional theory-based MD (DFTMD) simulations Car and Parrinello (1985); Marx and Hutter (2000); Car (2002); Kuo and Mundy (2004); Kühne et al. (2007, 2011); Khaliullin and Kühne (2013) also offer a predictive platform for understanding density profiles and determining the thickness of the air-liquid interfaces Blas et al. (2001); Kühne et al. (2011); Kessler et al. (2015); Pezzotti, Galimberti, and Gaigeot (2017).
Inspired by the above experimental and simulation results, and with the motivation of capturing realistic HB dynamics at interfaces, we have designed an approach based on two extreme scenarios of interface HB dynamics. In the first scenario, for the set of molecules located in the interface layer at a given sampling time, we use the Luzar-Chandler (LC) HB population operator Luzar and Chandler (1996a) to obtain the HB dynamics of these interface molecules. In the second scenario, taking inspiration from Luzar and Chandler’s HB population and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017), we have developed an interface HB (IHB) population operator. This operator aims to provide a refined understanding of HB dynamics specifically at the interface.
The Luzar-Chandler HB population is utilized to describe whether a pair of labeled molecules form H-bonds. And the characteristic function developed by Giberti and Hassanali describes whether a specific molecule belongs to the interface region. Therefore, our newly defined interface HB population can describe whether a labeled pair of molecules is within the interface and connected by H-bonds at any given moment. This dual condition provides a more detailed understanding of interfacial HB dynamics.
Due to the thermal motion of water molecules, two scenarios may occur. In the first scenario, the water molecules under observation might transition into the bulk phase. In the second scenario, if a water molecule resides in the interface region and its H-bonded partner moves outside the interface area, then such a pair of molecules will no longer be of concern. Based on the study of interfacial water dynamics by Liu et al. Liu, Harder, and Berne (2005), Gan et al. Gan et al. (2006a), Singh et al. Singh et al. (2013) and Jeon et al. Jeon et al. (2017), as well as the investigation into the time-dependent spectral evolution of H-bonded and free water molecules by Ojha and Kühne Ojha and Kühne (2021), our approach indicates that the HB dynamics derived from the first scenario will be slower the genuine interfacial HB dynamics. Conversely, the one obtained from the second scenario will exhibit a faster dynamics than the genuine one.
Building upon the foundation laid by methods reliant on the density criterion Townsend and Rice (1991); Beaglehole and Wilson (1993); Lie et al. (1993); Taylor, Dang, and Garrett (1996); Sokhan and Tildesley (1997); Alejandre, Tildesley, and Chapela (1995); Sokhan and Tildesley (1997); Morita and Hynes (2000); Paul and Chandra (2004); Morita (2006); Zhao et al. (2018), our approach introduces an alternative way of determining interface thickness through the analysis of the convergence of interfacial HB dynamics properties. This approach effectively bypasses the necessity of accounting for liquid density. As such, it offers another perspective for measuring the thickness of air-water interfaces. Furthermore, the principles underlying our approach hold potential for application to a broader range of systems, such as solution interfaces and ion shells, offering a flexible tool for interface studies.
II Methods
Due to molecular motions, the identity of molecules at the interface changes with time, and generally useful procedures for identifying interfaces must accommodate these motions. The air-water boundary is modeled with the Willard-Chandler instantaneous surface Willard and Chandler (2010); Pezzotti, Galimberti, and Gaigeot (2017); Serva et al. (2018). Figure  1 illustrates the obtained interfaces for one configuration of a slab of water.
Figure 1: A slab of water with the instantaneous surface $\mathbf{S}$ represented as a blue mesh on the upper and lower phase boundary. The grey surface, which represents an imaginary interface $\mathbf{S}^{\prime}$, is obtained by translating the surface to the inside of the system along the normal ($z$-axis) by distance $d$. This distance $d$ is then utilized to measure the thickness of the air-water interface.
For the slab in the cuboid simulation box, an imaginary surface ${\mathbf{S}}^{\prime}(t)$ is obtained by translating the surface ${\mathbf{S}}(t)$ along the system’s normal (into bulk) to a distance $d$. The region between the two surfaces ${\mathbf{S}}(t)$ and ${\mathbf{S}}^{\prime}(t)$ is defined as the air-water interface. Below we will combine two extreme scenarios to investigate the HB dynamics at the instantaneous air-water interface.
II.1 Scenario 1: HB dynamics based on the Luzar-Chandler HB population $h[{\bf r}(t)]$
As the first scenario we use the Luzar-Chandler HB population operator Luzar and Chandler (1996a) and employ a technique that samples water molecules right at the instantaneous interface for certain sampling time points. During the $t_{\rm traj}$ period, most water molecules only undergo thermal motion near their equilibrium position. However, a small number of molecules that are at the interface at time $t$ may enter the bulk phase at a later time. Due to the inclusion of H-bonds in the bulk phase, this scenario tends to underestimate the breaking rate of the H-bonds at the interface. This Luzar-Chander HB population-based method includes three steps as follows.
1. Sub-trajectories: Sub-trajectories with a specific length of time, $t_{\text{traj}}\sim 10$ ps are selected. Here, $t_{\text{traj}}$ is long enough to observe HB dynamics but short enough that the molecules do not complete their transition across the interface Madarasz, Rossky, and Turi (2007).
2. Sampling: At multiple, evenly spaced moments within $t_{\text{traj}}$, we identify a layer at the interface of a specified thickness $d$ (see Fig. 1) and select the water molecules within this layer for analysis. For each of these moments, the Luzar-Chandler HB population-based correlation functions Luzar and Chandler (1996a) for these molecules across the sub-trajectory $t_{\text{traj}}$ are calculated.
3. Statistics: The average properties over all sampled moments and all sub-trajectories are calculated.
For details on this LC HB operator-based method (LC method), see Appendix C.
II.2 Scenario 2: HB dynamics based on interface HB population $h^{(\text{s})}[{\mathbf{r}}(t)]$
To capture the other extreme of interfacial HB dynamics, after determining the instantaneous interface, we introduce an interface HB population operator $h^{(\text{s})}[{\mathbf{r}}(t)]$ as follows: It has a value of 1 when a tagged molecular pair $i,j$ are H-bonded and both molecules are at the interface with a thickness $d$, and 0 otherwise:
$h^{(\text{s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{j\in\mathcal{S}_{i}(t)};\\ 0&&{\text{otherwise,}}\end{array}\right.$ (4)
where ${\mathcal{S}_{i}(t)}$ denotes the set of water molecules that are H-bonded with molecule $i$ at time $t$, and $\mathcal{I}(d;t)$ is the instantaneous interface layer with thickness $d$ at time $t$, as defined previously. The definition of $h^{(\text{s})}$ combines the Luzar-Chandler’s HB population Luzar and Chandler (1993a, 1996a) $h$ and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017). Then the correlation function $c^{(\text{s})}(t)$ that describes the fluctuation of H-bonds at the interface:
$c^{(\text{s})}(t)=\langle h^{(\text{s})}(0)h^{(\text{s})}(t)\rangle/\langle h^{(\text{s})}\rangle,$ (5)
can be obtained.
Similar to functions $n(t)$ and $k(t)$ in Ref. 76 (Eq.s  12 and  13 in Appendix A), the corresponding correlation function
$n^{(\text{s})}(t)=\langle h^{(\text{s})}(0)[1-h^{(\text{s})}(t)]h^{(\text{d,s})}\rangle/\langle h^{(\text{s})}\rangle,$ (6)
and interface reactive flux function
$k^{(\text{s})}(t)=-{dc^{(\text{s})}(t)}/{dt}$ (7)
are obtained. The $h^{(\text{d,s})}(t)$ is 1 when a tagged pair of water molecules $i$, $j$ is at the interface and the inter-oxygen distance between the two molecules is less than the cutoff radius $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise, i.e.,
$h^{(\text{d,s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{|\mathbf{r}_{i}-\mathbf{r}_{j}|<r_{\text{OO}}^{\text{c}};}\\ 0&&{\text{otherwise.}}\end{array}\right.$ (11)
Therefore, $n^{(\text{s})}(t)$ represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules at the interface are unbonded but remain at the interface and separated by less than $r_{\text{OO}}^{\text{c}}$; $k^{(\text{s})}(t)$ measures the effective decay rate of H-bonds at the interface. The functions defined in Eq.s 57 are used to determine the reaction rate constants of breaking and reforming and the lifetimes of H-bonds at the interface by Luzar and Chandler (1996a); Khaliullin and Kühne (2013)
$k^{\rm(s)}(t)=kc^{\rm(s)}(t)-k^{\prime}n^{\rm(s)}(t).$
In this IHB scenario, choosing the water molecules and H-bonds at the interface is accurate. However, for some special H-bonds, if it connects such two water molecules, one is at the interface and the other is in bulk water, the HB breaking reaction rate of such H-bonds will be increased. Therefore, in contrast to the LC method used in scenario 1, the IHB method used in this scenario overestimates the HB breaking rate constant.
The actual HB dynamics at the interface are expected to lie between the results obtained by the LC method and the IHB method. Consequently, by integrating these two scenarios, we can achieve a more precise characterization of HB dynamics at the interface.
III Results and Discussions
In this section, we implement our two-extreme approach to two distinct system properties: HB population auto-correlation functions and HB reaction rate constants. We then determine the thickness of the air-water interface based on each of these properties.
III.1 HB population auto-correlation functions
Figure 2: The $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ obtained from the ADH criterion of HB definition for (a) scenario 1 (LC); (b) scenario 2 (IHB). Two features can be found: (i). As $d$ increases, both $c(t)$ and $c^{\rm(s)}(t)$ eventually approach a stable function. (ii). The decay rate of $c^{(\text{s})}(t)$ in scenario 2 surpasses $c(t)$ in scenario 1. This is visually represented by two dashed directed line segments, positioned identically to the graphs.
Figure 2 illustrates the dynamic evolution of $c(t)$ and $c^{\rm(s)}(t)$ for various values of $d$. When $d$ is equal to or greater than 4.0 (Å), the time evolution curves of $c(t)$ and $c^{\rm(s)}(t)$ become largely invariant to further increases in $d$. This shows that one can calculate the thickness of the interface under different conditions with the help of the dependence of the correlation function on $d$. Analyzing both scenarios, it is evident that the decay rate of the correlation function $c^{(\text{s})}(t)$ in scenario 2 (IHB) surpasses that in scenario 1 (LC).
Figure  3 shows the $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps). It shows the $d$-dependence clearly from another angle: we selected the three reference time intervals on the $t$-axis in Fig. 2 and recorded the values of the correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, at these time intervals for each $d$. Comparing $c(t)$ and $c^{(\text{s})}(t)$ in Fig. 3, $c(t)$ is always slightly larger than $c^{(\text{s})}(t)$ for the same $d$.
In Fig. 3, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ increase as $d$ increases, and eventually their change rates approach 0 when $d$ is large enough. Since $c(t)$ and $c^{(\text{s})}(t)$ are the upper and lower bounds of the real interface correlation functions $c^{r}(t^{*})$ respectively. It logically follows that the $d$-dependence of $c^{r}(t^{*})$ has the same changing trend. Then we have the equations ${\rm d}C/{\rm d}d=0$, for $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, with respect to $d$. For $C=c(t^{*})$ and $C=c^{\rm(s)}(t^{*})$, we can find solutions $d_{f1}$ and $d_{f2}$, respectively. Here $d_{f1}$ and $d_{f2}$ are the thickness of the air-water interface obtained from the LC method and the IHB method, respectively. Then $d_{f}=(d_{f1}+d_{f2})/2$ is obtained as the thickness of the interface for a given $t^{*}$.
Figure 3: The $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps), revealing key insights into the dynamics of the air-water interface: (i). As $d$ increases, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ show an upward trend, with their rates of change gradually approaching 0. (ii). For each $t^{*}$, $c(t)$ is always slightly larger than $c^{(s)}(t)$ for the same $d$. (iii). The thickness $d_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, obtained by the LC method (dot-dashed lines) and the IHB method (solid lines), respectively.
The correlations $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ respectively describe the relaxation characteristics of HB dynamics within the interface. For both the LC and IHB methods, the values of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ no longer change significantly with the increase of $d$, when $d$ increases to 4.0 (Å), ie., $d_{f1}\approx d_{f2}=4.0$ (Å). Therefore, we obtain the thickness of the air-water interface in the simulations, i.e., $d_{f}=4.0$ (Å). Note that the ADH criterion of H-bonds is used here. Such an observation is in harmony with the findings based on the AHD criterion, as elaborated in Appendix A.
III.2 HB reaction rate constants
Figure 4: The breaking HB reaction rate constants $k_{\rm LC}$ and $k_{\rm IHB}$ obtained by the IHB and LC methods respectively. (i). Both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically to the HB breaking rate $k_{\rm bulk}$ for the bulk water as $d$ increases. (ii). $k_{\rm LC}$ is always smaller than $k_{\rm IHB}$ for the same $d$. (iii). The thickness $d^{\prime}_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of the rate constant $k_{\rm IHB}$ and $k_{\rm LC}$, obtained by the IHB method and the LC method, respectively.
We further examined how the reaction rate constants of H-bonds at the air-water interface vary with $d$. For further details regarding the HB reaction rate constants, please refer to Appendix B.
In Fig. 4, we compare the breaking HB reaction rate constantsLuzar and Chandler (1996a); Luzar (2000), $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the IHB and LC methods, respectively. We found that both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically as $d$ increases. When $d$ is large, both rate constants $k_{\rm IHB}$ and $k_{\rm LC}$ also no longer change with $d$. In addition, the HB breaking rate $k_{\rm bulk}$ for the bulk water is also drawn as a reference. As $d$ increases, both $k_{\rm LC}$ and $k_{\rm IHB}$ gradually close to the rates $k_{\rm bulk}$. It also shows that the HB beaking rate constant $k_{\rm IHB}$ is relatively larger than $k_{\rm LC}$. This difference is related to the definitions of HB populations, $h(t)$ and $h^{\text{(s)}}(t)$. The definition of $h^{(\text{s})}(t)$ makes the HB break rate at the interface increase. The LC method, which is based on Luzar-Chandler HB population $h(t)$, retains the original rate constant of H-bonds, but it may include the contribution of bulk water molecules to the rate constant.
Similar to the HB auto-correlation functions, we have the equations ${\rm d}K/{\rm d}d=0$, for $k_{\rm IHB}$ and $k_{\rm LC}$, with respect to $d$. For $K=k_{\rm IHB}$ and $K=k_{\rm LC}$, we can find solutions $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$, respectively. Here $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$ are also the HB reaction rate-based thickness of the interface obtained from the LC and IHB method, respectively. Then we obtain the thickness of the real air-water interface as $d^{\prime}_{f}=(d^{\prime}_{f1},d^{\prime}_{f2})/2\approx 4.0$ (Å). This result is supported by our calculation of the $d$-dependence of the HB reforming rate constants, ie., $k^{\prime}_{\rm LC}$ and $k^{\prime}_{\rm IHB}$. For more details, please refer to Fig. B7 in Appendix B).
III.3 Orientational relaxation of the OH stretch and other supports
To verify our conclusion on the thickness of the air-water interface from the interfacial HB dynamics, we calculated the orientational relaxation of the OH stretch from the perspective of anisotropy decay. As shown in Fig. 5, the values of the orientational correlation function $C_{2}(t)$ at four reference times $t^{*}=1,2,5$ (ps) are plotted. It can be seen that when the interface thickness $d$ is small, the OH orientational relaxation time $\tau_{2}$ of the $C_{2}(t)$ increases linearly with $d$. The $\tau_{2}$ for the neat water surface is $1.4$ to $2.0$ (ps), with std. error $\delta d=0.1$ (ps), which is consistent with the vibrational relaxation time of the free OH for the neat water ichi Inoue et al. (2020).
Similar to the HB population correlation functions $c(t)$ and $c^{\rm(s)}(t)$ and the HB reaction rates, $k_{\rm LC}$ and $k^{\prime}_{\rm LC}$, and $k_{\rm IHB}$ and $k^{\prime}_{\rm IHB}$, as $d$ increases, $C_{2}(t^{*})$ and $\tau_{2}$ converges to a fixed value, which characterizes the decay time of the orientation relaxation process of OH bonds in bulk water. From the convergence trend of $\tau_{2}$ in Fig. 5(b), we find that at the interface with a thickness greater than 4.0 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water. Therefore, we arrive at a consistent conclusion on the issue of estimating the thickness of the air-water interface, from the perspective of HB dynamics and OH reorientation relaxation. Furthermore, the interface thickness aligns with values obtained through experimental measurements and AIMD simulations, as detailed in Table 1.
Figure 5: The $d$-dependence of (a) $C_{2}(t^{*})$ and (b) $\tau_{2}(t^{*})$ for water molecules at the air-water interface with $t^{*}=1,2,5$ (ps), respectively. As $d$ increases to $d=4$ (Å), $C_{2}(t^{*})$ and $\tau_{2}$ converge to a fixed value, respectively.
Table 1: The air-water interface thickness obtained by experiments and computer simulations ("-" denotes without using "10–90" thickness).
Method Thickness Notation $T$ (K) $d$ (Å)
Ellipsometry (Rayleigh) - 293.15 3.0
Ellipsometry (Raman and Ramdas (1927)) - 293 5.0
Ellipsometry (McBain, Bacon, and Bruce (1939)) - 293 $\geq$2.26
Ellipsometry (Kinosita and Yokota (1965)) - 293 7.1
X-Ray Reflectivity(Braslau et al. (1988)) - 298 3.24$\pm$ 0.05
MD (TIP4P/FQ, Liu et alLiu, Harder, and Berne (2005)) 10–90 298.15 3.5
MD (TIP4P-POL2, Kuo et alKuo et al. (2006)) 10–90 298 3.9
MD (SPC-FW, Wick et alWick et al. (2007)) 10–90 298 3.2
MD (D-C, Wick et alWick et al. (2007)) 10–90 298 3.2
MD (TIP4P, Wick et alWick et al. (2007)) 10–90 298 3.4
MD (SPC/E, Fan et alFan et al. (2009)) 10–90 300 6.0
MD ( SPC/E, Willard and ChandlerWillard and Chandler (2010)) 10–90 298 8.0
MD (SPC/E, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.27
MD (TIP4P/2005, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.05
CPMD/BLYP (Kuo et alKuo et al. (2006)) 10–90 300 1.7
BOMD/BLYP-D (Kühne et al. Kühne et al. (2011)) 10–90 300 3.8
BOMD/BLYP-D2 (Pezzotti et alPezzotti, Galimberti, and Gaigeot (2017) ) 10–90 315 3.5
BOMD/BLYP-D3 (Jeon el al. Jeon et al. (2017)) 10–90 298 $\sim$3.0
BOMD/BLYP-D3/LC&IHB (this work) - 300 4.0
IV Conclusions
In this study, we have developed a "two-extremes" approach to investigate the HB dynamics at the air-water interface and to determine the interface’s thickness. One extreme scenario underestimates the HB breaking rate constant, while the other overestimates it, implying that each scenario provides only a partial insight into the interfacial HB dynamics. Subsequently, based on the DFTMD simulations, we have applied our approach to two distinct system properties: HB relaxation and HB reaction rate constants at the air-water interface. Our results across all these properties indicate that the predictions from both extreme scenarios converge as the thickness of the air-water interface increases to 4.0 Å. Thus, we have reason to believe the thickness, which falls between these two extremes, converges at this critical value.
This work introduces an approach to complement existing ones for investigating the air-water interface from a fresh perspective. Through HB dynamics of the air-water interface, interfacial properties, such as thickness in this case, can be obtained through a method analogous to the squeeze theorem. Beyond the scope of HB dynamics, this approach can be extended to other properties like molecular orientation distribution Rao, Tao, and Wang (2003); Gan et al. (2006b) and SFG spectrum Nihonyanagi et al. (2011, 2015); Tang et al. (2020), as well as other solution interfacial surfaces where statistical properties of the interface and bulk phase differ significantly. Looking ahead, our study could also inspire researchers to examine the hydration shells of ions, and then study the effects of ions through HB dynamics.
Acknowledgements.
This work was supported by the CSC | Chinese Government Scholarship (No. 2011113214). We thank Marialore Sulpizi for the useful discussion. The simulations were performed on the Mogon ZDV cluster in Mainz and on the Cray XE6 (Hermit) at the HRLS supercomputing center in Stuttgart.
Appendix A Hydrogen bond correlation functions
[…]
We use a configuration $r(t)$ to denote the positions of all the atoms in the system at time $t$. Either of the criteria above allows one to define an HB population $h[r(t)]=h(t)$, which equals 1 when a tagged pair of molecules are H-bonded, and 0 otherwise. The fluctuation in $h(t)$ from its time-independent equilibrium average is defined by Chandler (1987) $\delta h=h(t)-\langle h\rangle$. The probability that a specific tagged pair of molecules is H-bonded in a large system is extremely small [^1] The average number of H-bonds in an equilibrium of $N$ water molecules is: $\frac{1}{2}N(N-1)\langle h\rangle$, i.e., $\frac{1}{2}N(N-1)\langle h\rangle\sim N$. It implies that $\frac{1}{2}N\langle h\rangle\sim 1$, or $\langle h\rangle\sim\frac{2}{N}$. Therefore, for a large system, $\langle h\rangle$ is extremely small., then $\delta h(t)=h(t)$. Therefore, the correlation of $\delta h(t)$ can be written as
$\langle\delta h(0)\delta h(t)\rangle=\langle h(0)h(t)\rangle,$
[…]
The correlation function $c(t)=\langle h(0)h(t)\rangle/\langle h\rangle$ describes the structural relaxation of H-bonds Chandra (2000); Benjamin (2005). Here the average $\langle h\rangle$ of the HB population is the probability that a pair of randomly chosen water molecules in the system is H-bonded at any time $t$. The function $c(t)$ measures correlation in $h(t)$ independent of any possible bond-breaking events, and it relaxes to zero when $t$ is large Rapaport (1983).
Because the thermal motion can cause distortions of H-bonds from the perfectly tetrahedral configuration, water molecules show a librational motion on a time scale of $\sim$ 0.1 ps superimposed to rotational and diffusional motions ($>1$ ps), which causes a time variation in interaction parameters. A new HB population $h^{\text{(d)}}(t)$ was also defined to obviate the distortion of real HB dynamics due to the above geometric definition Sciortino and Fornili (1989); Chandra (2000). It is 1 when the inter-oxygen distance of a particular tagged pair of water molecules is less than $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise. The H-bonds between a tagged molecular pair that satisfy the condition $h^{\text{(d)}}(t)=1$ may have been broken, but they may more easily form H-bonds again. The correlation function
$n(t)=\langle h(0)[1-h(t)]h^{\text{(d)}}(t)\rangle/\langle h\rangle$ (12)
represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules are unbonded but remain separated by less than $r_{\mathrm{OO}}^{\mathrm{c}}$ Chandra (2000).
The rate of HB relaxation to equilibrium is characterized by the reactive flux Luzar (2000)
$k(t)=-\frac{\mathrm{d}c(t)}{\mathrm{d}t},$ (13)
which quantifies the rate that an initially present HB breaks at time $t$, independent of possible breaking and reforming events in the interval from 0 to $t$. Therefore, $k(t)$ measures the effective decay rate of an the initial set of H-bonds Chandler (1987). For bulk water, a $\sim 0.2$-ps transient period exists, during which $k(t)$ changes quickly from its initial value Starr, Nielsen, and Sta (2000). However, at longer times, $k(t)$ is independent of the HB definitions.
The dynamic evolution of HB population correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, in the two scenarios, is shown in Fig.  A6a and b, respectively. This figure applies the AHD criteria to H-bonds. A critical thickness $d_{f}$ (equal to or greater than 4.0 (Å)) exists. Beyond this thickness, the time evolution curves of the correlation function $c(t)$ and $c^{\rm(s)}(t)$ do not significantly change with further increases in $d$.
A comparison between the two scenarios shows that in scenario 2, the correlation function $c^{\rm(s)}(t)$ decays faster than in scenario 1. This is visually indicated by a black directed line segment in the figure. This observation is consistent with conclusions drawn from another criterion, the ADH criterion, as we have shown in the main text.
The result, where $c^{\rm(s)}(t)$ decays faster in scenario 2 compared to $c(t)$ in scenario 1, aligns with findings from the ADH standard. This suggests a level of agreement between the different criteria applied to analyze the air-water interface.
Figure A6: The $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ obtained from the AHD criteria: (a) LC; (b) IHB. In scenario 1, the sampling is performed every 4 ps.
Appendix B Hydrogen bond breaking and reforming rate constants
Assume that each HB acts independently of other H-bonds Luzar and Chandler (1996b); Luzar (2000), and due to the detailed balance condition, one obtains $\tau_{\mathrm{HB}}={(1-\langle h\rangle)}/{k}$, where $k$ is the rate constant of breaking an HB, i.e., the forward rate constant Chandler (1986, 1978). Correspondingly, the backward rate constant $k^{\prime}$ is represented by the rate constant from the HB-on state to the HB-off state for a tagged pair of molecules. Based on the functions (see Appendix  A) $n(t)$, $h(t)$, $h^{\text{(d)}}(t)$, and $k(t)$, Khaliullin and Kühne (2013) have obtained the ratio $k/k^{\prime}$ of HB breaking and reforming rate constants in bulk water, and then the lifetime and relaxation time of H-bonds from simulations. Here, for the air-water interface, we obtain the optimal solution range of $k$ and $k^{\prime}$ from the relationship between the reactive flux $k(t)$ and the correlation functions $c(t)$ and $n(t)$:
$k(t)=kc(t)-k^{\prime}n(t).$ (14)
We obtain the optimal value of the rate constants, $k$ and $k^{\prime}$, by the least squares fit of $k(t)$, $c(t)$, and $n(t)$ beyond the transition phase. The function $c(t)$ is regarded as a column vector composed of $(c_{1},\cdots,c_{P})^{\mathrm{T}}$, and is denoted as ${\bf c}$, with $c_{i}$ representing the value of correlation $c(t)$ at $t=i$. Similarly, $n(t)$ and $k(t)$ can also be denoted as ${\bf n}$ and ${\bf k}$, respectively. Then, the rate constants $k$ and $k^{\prime}$ are determined from the matrix ${\bf A}=\begin{bmatrix}{\bf c}&{\bf n}\end{bmatrix}$:
$\begin{bmatrix}k\\ -k^{\prime}\end{bmatrix}=({\bf A}^{\mathrm{T}}{\bf A})^{-1}{\bf A}^{\mathrm{T}}{\bf k}.$ (15)
For bulk water and the air-water interface, the optimal $k$ and $k^{\prime}$ are reported in Table  B2 and  B3.
Table B2: The rate constants $k$ and $k^{\prime}$ for the bulk water and the air-water interface (the time region $0.2<t<$ 2 (ps)).
Criterion $k$ (b) [^2] The unit for $k$ ($k^{\prime}$) is ps-1. b: bulk; i: interface. $k^{\prime}$ (b) $\tau_{\mathrm{HB}}$ (b) [^3] The unit for $\tau_{\mathrm{HB}}=1/k$) is ps. $k$ (i) $k^{\prime}$ (i) $\tau_{\mathrm{HB}}$ (i)
ADH 0.296 0.988 3.380 0.323 0.765 3.101
AHD 0.288 1.149 3.470 0.314 0.887 3.184
Table B3: The $k$ and $k^{\prime}$ for the bulk water and the air-water interface (the time region $122<t<$ 12 (ps)).
Criterion $k$ (b) $k^{\prime}$ (b) $\tau_{\mathrm{HB}}$ (b) $k$ (i) $k^{\prime}$ (i) $\tau_{\mathrm{HB}}$ (i)
ADH 0.115 0.039 8.718 0.157 0.068 6.372
AHD 0.105 0.047 9.496 0.155 0.088 6.472
Figure B7: The $d$-dependence of the rate constant $k^{\prime}_{\rm IHB}$ and $k^{\prime}_{\rm LC}$, obtained by the IHB (solid lines) and the Luzar-Chandler (dot-dashed lines) methods, respectively. The corresponding reaction rate $k^{\prime}_{\rm bulk}$ for bulk water is also represented as dashed lines. The ADH criterion of H-bonds is used and the fits are carried on the time region $120.2<t<$ 12 (ps).
To obtain $k$ and $k^{\prime}$, we perform the fitting in short and long-time regions, respectively. Since HB dynamics show different relaxation behaviors in the time of $\sim 1$ ps and in that greater than $\sim 1$ psStarr, Nielsen, and Stanley (2000), we will fit the short and long time intervals respectively. We note that in the long time region ($122<t<$ 12 (ps)), the value of HB lifetime $\tau_{\mathrm{HB}}$ is larger than that in short one ($20.2<t<$ 2 (ps)), no matter for the bulk water or the air-water interface. A larger $\tau_{\mathrm{HB}}$ value means that the distance between a pair of water molecules stays within $r_{\mathrm{OO}}^{c}$ for a longer time.
Appendix C Details of Scenario 1: the LC method
In the LC scenario, to obtain interface HB dynamics, we use molecule sampling at the instantaneous interface. Let $T_{\text{e}}$ be the time it takes for all water molecules in the simulation box to traverse the interface and bulk phase. If the trajectory length $t_{\text{traj}}$ satisfies the condition $\tau_{\text{HB}}\ll t_{\text{traj}}\ll T_{\text{e}}$, we sample time points $t_{i}$ $(i=1,\cdots,M)$ that are evenly spaced on the trajectory, then obtain interface HB dynamics using the following procedure:
  1. 1.
    For each sampling time $t_{i}$, define an interface with a thickness $d$ (see Fig. 1). Then, select a set $S_{i}$ of water molecules in the interfacial layer $\mathcal{I}(d;t)$, i.e., $\mathbf{r}(t_{i})\in\mathcal{I}(d;t)$, and calculate the correlation functions $c(t)$, $n(t)$ and $k(t)$ through $t_{\text{traj}}$ for these water molecules.
  2. 2.
    Calculate average functions of the correlation functions $c(t)$, $n(t)$, and $k(t)$ respectively over all sampling time points and all trajectories.
  3. 3.
    Calculate reaction rate constants of breaking and reforming in the interfacial layer $\mathcal{I}(d;t)$ by Eq. 14.
In scenario 1, the sampling is performed every 4 ps.
Appendix D Rotational anisotropy decay of OH stretch at the water/vapor interface
We examine the results for the water interface thickness from the interfacial HB dynamics by polarization anisotropy decay of OH stretch. The polarization anisotropy decay of the OH stretch can also provide information on the dynamics of water  Woutersen and Bakker (1999); Cowan et al. (2005); Kraemer et al. (2008); Yagasaki, Ono, and Saito (2009). We assume that the anisotropy decay $R(t)$ is only due to the orientational relaxation of water molecules. Within the Condon approximation  Schmidt, Corcelli, and Skinner (2005), it is directly related to the orientational correlation function $C_{2}(t)$ Tokmakoff (1996); Rezus and Bakker (2006); Yagasaki, Ono, and Saito (2009); Bakker and Skinner (2010) through $R(t)=\frac{2}{5}C_{2}(t)$. The $C_{2}(t)$ is given by the rotational time-correlation function
$C_{2}(t)=\langle P_{2}(\hat{u}(0)\cdot\hat{u}(t))\rangle,$ (16)
[…]
For the air-water interface, using the method in section II, we first obtain the instantaneous interface with thickness $d=1,\cdots,6$ (Å). Then, for each value of $d$, the LC method is used to obtain $C_{2}(t)$ for water molecules at the air-water interface.
Here we introduce a fitting with an exponential and provide a decay time $\tau_{2}$, i.e., $C_{2}(t)=Ae^{-t/\tau_{2}}$, with $A=1.0$ for different thickness values, and the $d$-dependence of $C_{2}(t^{*})$ and $\tau_{2}(t^{*})$ is shown in Fig. 5. When the interface thickness $d$ is smaller than 4.0 (Å), the decay time $\tau_{2}$ increases linearly with $d$. The OH orientational relaxation time $\tau_{2}$ for the neat water surface is from $1.4$ to $2.0$ (ps), with std. error $\delta d=0.1$ (ps), which is consistent with the vibrational relaxation time of the free OH for the neat water ichi Inoue et al. (2020).
Similar to the correlation functions, $c(t)$ and $k(t)$, for interfacial water molecules, as $d$ increases, $\tau_{2}$ converges to a fixed value, which characterizes the decay time of the orientation relaxation process of OH bonds in bulk water. From the convergence trend of $\tau_{2}$ in Fig. 5(b), we find that at the interface with a thickness greater than 4.0 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water.
Appendix E Computational Methods
To describe the subtleties of H-bonding in water Laasonen et al. (1993), we have performed DFTMD simulations Marx and Hutter (2000) for bulk water and the air-water interface. The simulations make use of technologies that have been successfully tested on water and solutions Khatib et al. (2016a, b); Khatib and Sulpizi (2017), namely the Goedecker-Teter-Hutter (GTH) pseudopotentials Goedecker, Teter, and Hutter (1996); Hartwigsen, Goedecker, and Hutter (1998); Krack (2005), Generalised Gradient Approximation (GGA) of the exchange-correlation functional Becke (1988); Lee, Yang, and Parr (1988), and dispersion force correction, DFT-D3 Grimme et al. (2010); Klimeš and Michaelides (2012). By eliminating the strongly bound core electrons, the GTH pseudopotentials reduce the number of occupied electronic orbitals that have to be treated in an electronic structure calculation. There are dual-space Gaussian-type pseudopotentials that are separable and satisfy a quadratic scaling to system size Lu et al. (2019). The GGA functionals generally describe the dipole and quadruple moments of the molecules quite well, and the DFT-D3 correction treats the van der Waals dispersion forces and improves the structural properties without more computational cost and thus can be used for the air-water interface.
Figure E1: The partial radial distribution functions (RDFs) for the simulated bulk water system.
Figure E2: Probability distribution $P(z)$, along the $z$-axis, of H2O in the slab.
[…]
The bulk water system consisted of 128 water molecules in a periodic box of size $15.64\times 15.64\times 15.64$3, and with a density of 1.00 g cm-3. The slab consisted of 128 water molecules in a periodic box of size $15.64\times 15.64\times 31.28$3. The length of each trajectory is $60$ (ps). The RDFs $g_{\mathrm{OO}}(r)$ and $g_{\mathrm{OH}}(r)$ for the bulk water systems are shown in Fig. E1. The probability distribution of O and H atoms in the simulated model of the air-water interface is shown in Fig. E2.

### v3

@@
 However, accurately determining the boundary of interfacial and bulk water, that is, the thickness of the air-water interface, still challenges experimentalists.
-Various simulation-based methods have been developed to estimate the thickness, converging on a range of approximately 3 to 10 Å.
-In this study, we introduce a novel approach, grounded in density functional theory-based molecular dynamics (DFTMD) simulations, to measure the air-water interface thickness, offering a different perspective based on prior research.
+Various simulation-based methods have been developed to estimate the thickness, converging on a range of approximately 3–10 (Å).
+In this study, we introduce a novel approach, grounded in density functional theory-based molecular dynamics and deep potential molecular dynamics simulations, to measure the air-water interface thickness, offering a different perspective based on prior research.
 To capture realistic HB dynamics in the air-water interface, two extreme scenarios of the interface HB dynamics are obtained: one underestimates the interface HB dynamics, while the other overestimates it.
-Surprisingly, our results suggest that the interface HB dynamics in both scenarios converge as the thickness of the air-water interface increases to 4.0 Å.
-This convergence point, indicative of the realistic interface thickness, is confirmed by our calculation of anisotropic decay of OH stretch at the air-water interface and supported by the results from recent ab initio molecular dynamics simulations.
+Surprisingly, our results suggest that the interface HB dynamics in both scenarios converges as the thickness of the air-water interface increases to 4 (Å).
+This convergence point, indicative of the realistic interface thickness, is also validated by our calculation of anisotropic decay of OH stretch and the free OH dynamics at the air-water interface.
 preprint: AIP/123-QED
@@
 (2006a) found that at the air-water interface, singly hydrogen (H)-bonded water molecules align almost parallel to the interface with limited orientational variation, while doubly H-bonded donor molecules orient their dipole vectors away from the liquid phase, highlighting diverse behaviors among interfacial water molecules.
-Almost concurrently, through time-resolved SFG vibrational spectroscopy, McGuire and ShenMcguire and Shen (2006) observed ultrafast vibrational dynamics at the interface, noting that the relaxation behaviors of interfacially bonded OH stretch modes on sub-picosecond timescales were akin to those in bulk water, encompassing spectral diffusion, vibrational relaxation, and thermalization.
+Almost concurrently, through time-resolved SFG vibrational spectroscopy, McGuire and ShenMcguire and Shen (2006) observed ultrafast vibrational dynamics at the interface, noting that the relaxation behaviors of interfacially bonded OH stretch modes on subpicosecond time scales were akin to those in bulk water, encompassing spectral diffusion, vibrational relaxation, and thermalization.
 Pioneering work by Tahara’s group Singh et al.
@@
 These collective insights enhance our understanding of the vibrational energy relaxation, HB dynamics, and interactions of water molecules at the air-water interface.
-Building upon this knowledge of interfacial behavior, significant efforts have also been directed towards quantifying the physical characteristics of the interface, including its thickness, which plays a crucial role in understanding its molecular interactions and behavior.
+Building upon this knowledge of interfacial behavior, significant efforts have also been directed toward quantifying the physical characteristics of the interface, including its thickness, which plays a crucial role in understanding its molecular interactions and behavior.
 The air-water interface thickness has been measured via ellipsometry Raman and Ramdas (1927); McBain, Bacon, and Bruce (1939); Kinosita and Yokota (1965), relative permittivity measurements Fumagalli et al.
@@
 (2020) to mention just a few.
-There is a consensus that the thickness of the air-water interface is about 3 to 10 Å Jungwirth and Tobias (2006); Sedlmeier, Horinek, and Netz (2009); Ishiyama and Morita (2009); Kuo et al.
+There is a consensus that the thickness of the air-water interface is about 3–10 (Å) Jungwirth and Tobias (2006); Sedlmeier, Horinek, and Netz (2009); Ishiyama and Morita (2009); Kuo et al.
 (2006); Wick et al.
@@
 (2015); Pezzotti, Galimberti, and Gaigeot (2017).
-Inspired by the above experimental and simulation results, and with the motivation of capturing realistic HB dynamics at interfaces, we have designed an approach based on two extreme scenarios of interface HB dynamics.
-In the first scenario, for the set of molecules located in the interface layer at a given sampling time, we use the Luzar-Chandler (LC) HB population operator Luzar and Chandler (1996a) to obtain the HB dynamics of these interface molecules.
+Nevertheless, DFTMD simulations are constrained by limitations in time and the number of molecules they can model.
+Traditional force field approaches, however, often lack the accuracy required to describe complex interface systems.
+Schran et al.
+(2021) Recently, deep potential molecular dynamics (DeePMD) simulations based on machine learning potential (MLPs) have emerged as a promising alternative, offering a solution to the accuracy-versus-efficiency dilemma in molecular simulations.
+Zeng et al.
+(2023) One of the most accurate MLPs for water is MB-pol, which accurately reproduces many properties of water across the phase diagram.
+Babin, Leforestier, and Paesani (2013); Babin, Medders, and Paesani (2014); Medders, Babin, and Paesani (2014); Bore and Paesani (2023); Muniz et al.
+(2021) Moreover, MB-pol has been used to obtain the VSFG and surface tension of air-water interface, demonstrating excellent agreement between theoretical predictions and experimental measurements.
+Medders and Paesani (2016)
+Inspired by the above experimental and simulation results, and with the motivation of capturing realistic HB dynamics at interfaces, we have designed an approach based on two extreme scenarios of interface HB dynamics by utilizing the trajectories of DFTMD and DeePMD simulations based on MB-pol.
+In the first scenario, for the set of molecules located in the interface layer at given sampling times, we use the Luzar-Chandler (LC) HB population operator Luzar and Chandler (1996) to obtain the HB dynamics of these interface molecules.
 In the second scenario, taking inspiration from Luzar and Chandler’s HB population and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017), we have developed an interface HB (IHB) population operator.
@@
 (2017), as well as the investigation into the time-dependent spectral evolution of H-bonded and free water molecules by Ojha and Kühne Ojha and Kühne (2021), our approach indicates that the HB dynamics derived from the first scenario will be slower the genuine interfacial HB dynamics.
-Conversely, the one obtained from the second scenario will exhibit a faster dynamics than the genuine one.
+Conversely, the one obtained from the second scenario will exhibit faster dynamics than the genuine one.
 Building upon the foundation laid by methods reliant on the density criterion Townsend and Rice (1991); Beaglehole and Wilson (1993); Lie et al.
-(1993); Taylor, Dang, and Garrett (1996); Sokhan and Tildesley (1997); Alejandre, Tildesley, and Chapela (1995); Sokhan and Tildesley (1997); Morita and Hynes (2000); Paul and Chandra (2004); Morita (2006); Zhao et al.
+(1993); Taylor, Dang, and Garrett (1996); Sokhan and Tildesley (1997); Alejandre, Tildesley, and Chapela (1995); Morita and Hynes (2000); Paul and Chandra (2004); Morita (2006); Zhao et al.
 (2018), our approach introduces an alternative way of determining interface thickness through the analysis of the convergence of interfacial HB dynamics properties.
 This approach effectively bypasses the necessity of accounting for liquid density.
-As such, it offers another perspective for measuring the thickness of air-water interfaces.
+As such, it offers another perspective for measuring the thickness of the air-water interface.
 Furthermore, the principles underlying our approach hold potential for application to a broader range of systems, such as solution interfaces and ion shells, offering a flexible tool for interface studies.
@@
 Figure  1 illustrates the obtained interfaces for one configuration of a slab of water.
-Figure 1: A slab of water with the instantaneous surface $\mathbf{S}$ represented as a blue mesh on the upper and lower phase boundary.
-The grey surface, which represents an imaginary interface $\mathbf{S}^{\prime}$, is obtained by translating the surface to the inside of the system along the normal ($z$-axis) by distance $d$.
-This distance $d$ is then utilized to measure the thickness of the air-water interface.
-For the slab in the cuboid simulation box, an imaginary surface ${\mathbf{S}}^{\prime}(t)$ is obtained by translating the surface ${\mathbf{S}}(t)$ along the system’s normal (into bulk) to a distance $d$.
-The region between the two surfaces ${\mathbf{S}}(t)$ and ${\mathbf{S}}^{\prime}(t)$ is defined as the air-water interface.
+Figure 1: Slab of water containing 512 water molecules with the instantaneous surface $\mathcal{S}$ represented as a blue mesh on the upper and lower phase boundary.
+Grey surface, which represents an imaginary interface $\mathcal{S}^{\prime}$, is obtained by translating the surface to the inside of the system along the normal ($z$-axis) by distance $d$.
+This variable $d$ is then utilized to measure the thickness of the air-water interface.
+For the slab in the cuboid simulation box, an imaginary surface $\mathcal{S}^{\prime}(t)$ is obtained by translating the surface $\mathcal{S}(t)$ along the system’s normal (into bulk) to a distance $d$.
+The region between the two surfaces $\mathcal{S}(t)$ and $\mathcal{S}^{\prime}(t)$ is defined as the air-water interface.
 Below we will combine two extreme scenarios to investigate the HB dynamics at the instantaneous air-water interface.
-II.1 Scenario 1: HB dynamics based on the Luzar-Chandler HB population $h[{\bf r}(t)]$
-As the first scenario we use the Luzar-Chandler HB population operator Luzar and Chandler (1996a) and employ a technique that samples water molecules right at the instantaneous interface for certain sampling time points.
-During the $t_{\rm traj}$ period, most water molecules only undergo thermal motion near their equilibrium position.
-However, a small number of molecules that are at the interface at time $t$ may enter the bulk phase at a later time.
+II.1 Scenario 1: HB Dynamics Based on the Luzar-Chandler HB Population
+As the first scenario we use the Luzar-Chandler HB population and employ a technique that samples water molecules right at the instantaneous interface for certain sampling time points.
+In Scenario 1, we divide the simulation trajectory into multiple subtrajectories of length $t_{\rm traj}$.
+Within these subtrajectories, the majority of water molecules exhibit thermal fluctuations near their equilibrium positions.
+Meanwhile, a select group of molecules initially at the interface may transition into the bulk phase at a later time.
 Due to the inclusion of H-bonds in the bulk phase, this scenario tends to underestimate the breaking rate of the H-bonds at the interface.
-This Luzar-Chander HB population-based method includes three steps as follows.
-1.
-Sub-trajectories: Sub-trajectories with a specific length of time, $t_{\text{traj}}\sim 10$ ps are selected.
-Here, $t_{\text{traj}}$ is long enough to observe HB dynamics but short enough that the molecules do not complete their transition across the interface Madarasz, Rossky, and Turi (2007).
-2.
-Sampling: At multiple, evenly spaced moments within $t_{\text{traj}}$, we identify a layer at the interface of a specified thickness $d$ (see Fig.
-1) and select the water molecules within this layer for analysis.
-For each of these moments, the Luzar-Chandler HB population-based correlation functions Luzar and Chandler (1996a) for these molecules across the sub-trajectory $t_{\text{traj}}$ are calculated.
-3. Statistics: The average properties over all sampled moments and all sub-trajectories are calculated.
-For details on this LC HB operator-based method (LC method), see Appendix C.
-II.2 Scenario 2: HB dynamics based on interface HB population $h^{(\text{s})}[{\mathbf{r}}(t)]$
+In Section 3 of this paper, we will see this result, combined with the outcome from Scenario 2, can be used to estimate the thickness of the air-water interface.
+This method includes three steps as follows:
+(a) Sub-trajectories.
+Sub-trajectories with a specific length of time, $t_{\text{traj}}$, are selected.
+In this work, $t_{\rm traj}$ is set to be $40$ (ps), which is long enough to observe HB dynamics but short enough that not all the molecules complete their transition across the interface Madarász, Rossky, and Turi (2007).
+(b) Sampling.
+For each time step, we identify a pair of the air-water interfaces of a specified thickness $d$ as shown in Figure 1.
+At evenly spaced moments within $t_{\text{traj}}$, we select the water molecules within the interfaces.
+For the union of the interfacial water molecules picked at all these time moments, the Luzar-Chandler HB population-based correlation functions Luzar and Chandler (1996) across the sub-trajectory are calculated.
+(c) Statistics. The average correlation functions across all sub-trajectories are calculated.
+In this scenario, the computational procedure for calculating HB dynamics follows the same methodology as that used in works based on the Luzar-Chandler’s method (LC method)Luzar (2000); Benjamin (2005); Liu, Harder, and Berne (2005).
+For details on this method, please refer to the Supporting Information.
+II.2 Scenario 2: HB Dynamics Based on Interface HB Population
 To capture the other extreme of interfacial HB dynamics, after determining the instantaneous interface, we introduce an interface HB population operator $h^{(\text{s})}[{\mathbf{r}}(t)]$ as follows: It has a value of 1 when a tagged molecular pair $i,j$ are H-bonded and both molecules are at the interface with a thickness $d$, and 0 otherwise:
-$h^{(\text{s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{j\in\mathcal{S}_{i}(t)};\\ 0&&{\text{otherwise,}}\end{array}\right.$ (4)
-where ${\mathcal{S}_{i}(t)}$ denotes the set of water molecules that are H-bonded with molecule $i$ at time $t$, and $\mathcal{I}(d;t)$ is the instantaneous interface layer with thickness $d$ at time $t$, as defined previously.
-The definition of $h^{(\text{s})}$ combines the Luzar-Chandler’s HB population Luzar and Chandler (1993a, 1996a) $h$ and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017).
+$h^{(\text{s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{j\in\mathcal{B}_{i}(t)};\\ 0&&{\text{otherwise}}\end{array}\right.$ (4)
+where ${\mathbf{r}}(t)$ is the configuration of the system at time $t$, ${\mathbf{r}}_{i}$ is the position coordinate of the oxygen atom in the $i$-th water molecule, ${\mathcal{B}_{i}(t)}$ denotes the set of water molecules that are H-bonded with molecule $i$ at time $t$, and $\mathcal{I}(d;t)$ is the instantaneous interface layer with thickness $d$ at time $t$.
+The definition of $h^{(\text{s})}$ combines the Luzar-Chandler’s HB population Luzar and Chandler (1993, 1996) $h$ and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017).
 Then the correlation function $c^{(\text{s})}(t)$ that describes the fluctuation of H-bonds at the interface:
-$c^{(\text{s})}(t)=\langle h^{(\text{s})}(0)h^{(\text{s})}(t)\rangle/\langle h^{(\text{s})}\rangle,$ (5)
-can be obtained.
-Similar to functions $n(t)$ and $k(t)$ in Ref. 76 (Eq.s  12 and  13 in Appendix A), the corresponding correlation function
-$n^{(\text{s})}(t)=\langle h^{(\text{s})}(0)[1-h^{(\text{s})}(t)]h^{(\text{d,s})}\rangle/\langle h^{(\text{s})}\rangle,$ (6)
+$c^{(\text{s})}(t)=\frac{\langle h^{(\text{s})}(0)h^{(\text{s})}(t)\rangle}{\langle h^{(\text{s})}\rangle},$ (5)
+can be obtained. Similar to functions $n(t)$ and $k(t)$ in ref 84 (eqs 1 and 2 in Supporting Information), the corresponding correlation function
+$n^{(\text{s})}(t)=\frac{\langle h^{(\text{s})}(0)[1-h^{(\text{s})}(t)]h^{(\text{d,s})}\rangle}{\langle h^{(\text{s})}\rangle},$ (6)
 and interface reactive flux function
-$k^{(\text{s})}(t)=-{dc^{(\text{s})}(t)}/{dt}$ (7)
+$k^{(\text{s})}(t)=-\frac{\mathrm{d}c^{(\text{s})}(t)}{\mathrm{d}t}$ (7)
 are obtained.
-The $h^{(\text{d,s})}(t)$ is 1 when a tagged pair of water molecules $i$, $j$ is at the interface and the inter-oxygen distance between the two molecules is less than the cutoff radius $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise, i.e.,
+The $h^{(\text{d,s})}(t)$ is 1 when a tagged pair of water molecules $i$, $j$ is at the interface and the interoxygen distance between the two molecules is less than the cutoff radius $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise, i.e.,
 $h^{(\text{d,s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{|\mathbf{r}_{i}-\mathbf{r}_{j}|<r_{\text{OO}}^{\text{c}};}\\ 0&&{\text{otherwise.}}\end{array}\right.$ (11)
 Therefore, $n^{(\text{s})}(t)$ represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules at the interface are unbonded but remain at the interface and separated by less than $r_{\text{OO}}^{\text{c}}$; $k^{(\text{s})}(t)$ measures the effective decay rate of H-bonds at the interface.
-The functions defined in Eq.s 57 are used to determine the reaction rate constants of breaking and reforming and the lifetimes of H-bonds at the interface by Luzar and Chandler (1996a); Khaliullin and Kühne (2013)
+The functions defined in eqs 57 are used to determine the reaction rate constants of breaking and reforming and the lifetimes of H-bonds at the interface by Luzar and Chandler (1996); Khaliullin and Kühne (2013)
 $k^{\rm(s)}(t)=kc^{\rm(s)}(t)-k^{\prime}n^{\rm(s)}(t).$
 In this IHB scenario, choosing the water molecules and H-bonds at the interface is accurate.
-However, for some special H-bonds, if it connects such two water molecules, one is at the interface and the other is in bulk water, the HB breaking reaction rate of such H-bonds will be increased.
+However, for some special H-bonds, if it connects such two water molecules, one is at the interface and the other is in the bulk phase, the HB breaking reaction rate of such H-bonds will be increased.
 Therefore, in contrast to the LC method used in Scenario 1, the IHB method used in this scenario overestimates the HB breaking rate constant.
@@
 III Results and Discussions
-In this section, we implement our two-extreme approach to two distinct system properties: HB population auto-correlation functions and HB reaction rate constants.
+In this section, we apply our two-extreme approach to two system properties: HB population autocorrelation functions and HB reaction rate constants.
 We then determine the thickness of the air-water interface based on each of these properties.
-III.1 HB population auto-correlation functions
-Figure 2: The $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ obtained from the ADH criterion of HB definition for (a) scenario 1 (LC); (b) scenario 2 (IHB).
-Two features can be found: (i).
-As $d$ increases, both $c(t)$ and $c^{\rm(s)}(t)$ eventually approach a stable function.
-(ii).
-The decay rate of $c^{(\text{s})}(t)$ in scenario 2 surpasses $c(t)$ in scenario 1.
-This is visually represented by two dashed directed line segments, positioned identically to the graphs.
-Figure 2 illustrates the dynamic evolution of $c(t)$ and $c^{\rm(s)}(t)$ for various values of $d$.
-When $d$ is equal to or greater than 4.0 (Å), the time evolution curves of $c(t)$ and $c^{\rm(s)}(t)$ become largely invariant to further increases in $d$.
-This shows that one can calculate the thickness of the interface under different conditions with the help of the dependence of the correlation function on $d$.
-Analyzing both scenarios, it is evident that the decay rate of the correlation function $c^{(\text{s})}(t)$ in scenario 2 (IHB) surpasses that in scenario 1 (LC).
-Figure  3 shows the $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps).
-It shows the $d$-dependence clearly from another angle: we selected the three reference time intervals on the $t$-axis in Fig.
-2 and recorded the values of the correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, at these time intervals for each $d$.
-Comparing $c(t)$ and $c^{(\text{s})}(t)$ in Fig.
-3, $c(t)$ is always slightly larger than $c^{(\text{s})}(t)$ for the same $d$.
-In Fig.
-3, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ increase as $d$ increases, and eventually their change rates approach 0 when $d$ is large enough.
-Since $c(t)$ and $c^{(\text{s})}(t)$ are the upper and lower bounds of the real interface correlation functions $c^{r}(t^{*})$ respectively.
-It logically follows that the $d$-dependence of $c^{r}(t^{*})$ has the same changing trend.
-Then we have the equations ${\rm d}C/{\rm d}d=0$, for $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, with respect to $d$.
-For $C=c(t^{*})$ and $C=c^{\rm(s)}(t^{*})$, we can find solutions $d_{f1}$ and $d_{f2}$, respectively.
-Here $d_{f1}$ and $d_{f2}$ are the thickness of the air-water interface obtained from the LC method and the IHB method, respectively.
-Then $d_{f}=(d_{f1}+d_{f2})/2$ is obtained as the thickness of the interface for a given $t^{*}$.
-Figure 3: The $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps), revealing key insights into the dynamics of the air-water interface: (i).
-As $d$ increases, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ show an upward trend, with their rates of change gradually approaching 0.
-(ii).
-For each $t^{*}$, $c(t)$ is always slightly larger than $c^{(s)}(t)$ for the same $d$.
-(iii).
-The thickness $d_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, obtained by the LC method (dot-dashed lines) and the IHB method (solid lines), respectively.
-The correlations $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ respectively describe the relaxation characteristics of HB dynamics within the interface.
-For both the LC and IHB methods, the values of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ no longer change significantly with the increase of $d$, when $d$ increases to 4.0 (Å), ie., $d_{f1}\approx d_{f2}=4.0$ (Å).
-Therefore, we obtain the thickness of the air-water interface in the simulations, i.e., $d_{f}=4.0$ (Å).
-Note that the ADH criterion of H-bonds is used here.
-Such an observation is in harmony with the findings based on the AHD criterion, as elaborated in Appendix A.
+The computational details for DFTMD and DeePMD simulations are available in the Supporting Information.
+The results discussed in the following sections are derived from the trajectory data of the system containing 512 water molecules, modeled by MB-pol potential.
+The finite size effects of the main properties discussed in this article are in Supporting Information.
+Similar analyses for DFTMD simulations with 128 water molecules, and other system sizes in DeePMD simulations using MB-pol, are also provided in the Supporting Information.
+III.1 HB Population Autocorrelation Functions
+Figure 2: Autocorrelation functions $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ for (a) Scenario 1 (LC) and (b) Scenario 2 (IHB).
+Two notable features emerge: (i) As $d$ increases, both $c(t)$ and $c^{\rm(s)}(t)$ eventually approach a stable function.
+(ii) Decay rate of $c^{(\text{s})}(t)$ in Scenario 2 is greater than that of $c(t)$ in Scenario 1.
+This behavior is visually represented by two dashed directed line segments, positioned identically to the graphs.
+Functions $c(t)$ and $c^{(\text{s})}(t)$ on a log-log scale are also plotted in Supporting Information.
+Figure 2 illustrates the dynamic evolution of $c(t)$ and $c^{\rm(s)}(t)$ for various values of distance $d$.
+When $d$ is 4 (Å) or greater, $c(t)$ and $c^{\rm(s)}(t)$ become largely invariant to further increases in $d$.
+This indicates that one can determine the thickness of the interface under different conditions by examining how the correlation function depends on distance $d$.
+Analysis of both scenarios reveals that the decay rate of the correlation function $c^{(\text{s})}(t)$ in Scenario 2 (IHB) surpasses that in Scenario 1 (LC).
+Figure  3 displays the $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps).
+This figure provides a different perspective on $d$-dependence: by selecting three reference time intervals on the $t$-axis in Figure 2, the values of the correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, at these time intervals for each $d$ were recorded.
+Comparing $c(t)$ and $c^{(\text{s})}(t)$ in Figure 3, $c(t)$ in Scenario 1 is always slightly larger than $c^{(\text{s})}(t)$ in Scenario 2 for the same $d$.
+Figure 3: Dependence of the correlation functions on the distance $d$ at three reference time points $t^{*}=1,2,5$ (ps) provides key insights into the dynamics of the air-water interface: (i) As $d$ increases, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ exhibit an upward trend, with their rates of change gradually approaching 0.
+(ii) At each $t^{*}$, $c(t)$ is consistently slightly larger than $c^{\rm(s)}(t)$ for the same $d$.
+(iii) Air-water interface thickness $d_{f}=4$ (Å) is determined from the $d$-dependence of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, as calculated using the LC method (dot-dashed lines) and the IHB method (solid lines), respectively.
+In Figure 3, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ increase as $d$ increases, eventually reaching a point where their change rates approach 0 for sufficiently large values of $d$.
+Since $c(t)$ and $c^{(\text{s})}(t)$ serve as the upper and lower bounds of the real interface correlation function $c^{r}(t^{*})$ respectively, it logically follows that the $d$-dependence of $c^{r}(t^{*})$ exhibits the same trend.
+Consequently, we arrive at the equations ${\rm d}C/{\rm d}d=0$ for $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ with respect to $d$.
+Solving for $C=c(t^{*})$ and $C=c^{\rm(s)}(t^{*})$ yields solutions $d_{f1}$ and $d_{f2}$, respectively.
+Here $d_{f1}$ and $d_{f2}$ represent the thickness of the air-water interface as determined by the LC method and the IHB method, respectively.
+The average value $d_{f}=(d_{f1}+d_{f2})/2$ is then obtained as the thickness of the interface for a given $t^{*}$.
+The correlations $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ respectively describe the relaxation characteristics of HB dynamics within the air-water interface.
+For both the LC and IHB methods, the values of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ cease to change significantly when $d$ reaches 4 (Å), i.e., $d_{f1}\approx d_{f2}=4$ (Å).
+Thus, we determine the thickness of the air-water interface in the simulations to be $d_{f}=4$ (Å).
 III.2 HB Reaction Rate Constants
-Figure 4: The breaking HB reaction rate constants $k_{\rm LC}$ and $k_{\rm IHB}$ obtained by the IHB and LC methods respectively.
-(i).
-Both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically to the HB breaking rate $k_{\rm bulk}$ for the bulk water as $d$ increases.
-(ii).
-$k_{\rm LC}$ is always smaller than $k_{\rm IHB}$ for the same $d$.
-(iii).
-The thickness $d^{\prime}_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of the rate constant $k_{\rm IHB}$ and $k_{\rm LC}$, obtained by the IHB method and the LC method, respectively.
-We further examined how the reaction rate constants of H-bonds at the air-water interface vary with $d$. For further details regarding the HB reaction rate constants, please refer to Appendix B.
-In Fig.
-4, we compare the breaking HB reaction rate constantsLuzar and Chandler (1996a); Luzar (2000), $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the IHB and LC methods, respectively.
+We further examined how the reaction rate constants of H-bonds at the air-water interface vary with $d$.
+For further details regarding the HB reaction rate constants, please refer to Section 2 in Supporting Information.
+In Figure 4, we compare the breaking HB reaction rate constantsLuzar and Chandler (1996); Luzar (2000), $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the LC and IHB methods, respectively.
 We found that both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically as $d$ increases.
 When $d$ is large, both rate constants $k_{\rm IHB}$ and $k_{\rm LC}$ also no longer change with $d$.
-In addition, the HB breaking rate $k_{\rm bulk}$ for the bulk water is also drawn as a reference.
-As $d$ increases, both $k_{\rm LC}$ and $k_{\rm IHB}$ gradually close to the rates $k_{\rm bulk}$.
-It also shows that the HB beaking rate constant $k_{\rm IHB}$ is relatively larger than $k_{\rm LC}$.
+It also shows that the HB breaking rate constant $k_{\rm IHB}$ is relatively larger than $k_{\rm LC}$.
 This difference is related to the definitions of HB populations, $h(t)$ and $h^{\text{(s)}}(t)$.
-The definition of $h^{(\text{s})}(t)$ makes the HB break rate at the interface increase.
-The LC method, which is based on Luzar-Chandler HB population $h(t)$, retains the original rate constant of H-bonds, but it may include the contribution of bulk water molecules to the rate constant.
-Similar to the HB auto-correlation functions, we have the equations ${\rm d}K/{\rm d}d=0$, for $k_{\rm IHB}$ and $k_{\rm LC}$, with respect to $d$.
-For $K=k_{\rm IHB}$ and $K=k_{\rm LC}$, we can find solutions $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$, respectively.
-Here $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$ are also the HB reaction rate-based thickness of the interface obtained from the LC and IHB method, respectively.
-Then we obtain the thickness of the real air-water interface as $d^{\prime}_{f}=(d^{\prime}_{f1},d^{\prime}_{f2})/2\approx 4.0$ (Å).
-This result is supported by our calculation of the $d$-dependence of the HB reforming rate constants, ie., $k^{\prime}_{\rm LC}$ and $k^{\prime}_{\rm IHB}$.
-For more details, please refer to Fig.
-B7 in Appendix B).
+The definition of $h^{(\text{s})}(t)$ leads to an increased HB break rate at the interface.
+The LC method retains the original rate constant of H-bonds but may include contributions from bulk water molecules.
+Figure 4: Breaking HB reaction rate constants $k_{\rm LC}$ and $k_{\rm IHB}$ for the air-water interface simulated with 512 water molecules, obtained by the LC and IHB methods respectively.
+(i) Both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically to the HB breaking rate $k_{\rm bulk}$ for the bulk water as $d$ increases.
+(ii) $k_{\rm LC}$ is always smaller than $k_{\rm IHB}$ for the same $d$.
+(iii) Thickness $d^{\prime}_{f}=4$ (Å) of the air-water interface is obtained from the $d$-dependence of the rate constant $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the LC method and the IHB method, respectively.
+(Consistent results for systems with 125 and 216 water molecules, please refer to the Supporting Information.)
+Similar to the HB autocorrelation functions, we have the equations ${\rm d}K/{\rm d}d=0$, for $k_{\rm IHB}$ and $k_{\rm LC}$, with respect to $d$.
+For $K=k_{\rm IHB}$ and $K=k_{\rm LC}$, we identify solutions $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$, respectively.
+Here $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$ represent the HB reaction rate-based thickness of the interface obtained from the LC and IHB method, respectively.
+We then calculate the thickness of the real air-water interface as $d^{\prime}_{f}=(d^{\prime}_{f1}+d^{\prime}_{f2})/2\approx 4$ (Å).
+This result is supported by our calculation of the $d$-dependence of the HB reforming rate constants, namely, $k^{\prime}_{\rm LC}$ and $k^{\prime}_{\rm IHB}$.
+For more details, please refer to Figure S7 in Supporting Information.
 III.3 Orientational Relaxation of the OH Stretch and Other Supports
-To verify our conclusion on the thickness of the air-water interface from the interfacial HB dynamics, we calculated the orientational relaxation of the OH stretch from the perspective of anisotropy decay.
-As shown in Fig.
-5, the values of the orientational correlation function $C_{2}(t)$ at four reference times $t^{*}=1,2,5$ (ps) are plotted.
-It can be seen that when the interface thickness $d$ is small, the OH orientational relaxation time $\tau_{2}$ of the $C_{2}(t)$ increases linearly with $d$.
-The $\tau_{2}$ for the neat water surface is $1.4$ to $2.0$ (ps), with std.
-error $\delta d=0.1$ (ps), which is consistent with the vibrational relaxation time of the free OH for the neat water ichi Inoue et al.
+To verify our conclusion on the thickness of the air-water interface from the interfacial HB dynamics, we performed calculations on the orientational relaxation of the OH stretch from the perspective of anisotropy decayMoilanen et al.
 (2008).
-Similar to the HB population correlation functions $c(t)$ and $c^{\rm(s)}(t)$ and the HB reaction rates, $k_{\rm LC}$ and $k^{\prime}_{\rm LC}$, and $k_{\rm IHB}$ and $k^{\prime}_{\rm IHB}$, as $d$ increases, $C_{2}(t^{*})$ and $\tau_{2}$ converges to a fixed value, which characterizes the decay time of the orientation relaxation process of OH bonds in bulk water.
-From the convergence trend of $\tau_{2}$ in Fig.
-5(b), we find that at the interface with a thickness greater than 4.0 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water.
+Figure 5: Orientation correlation function $C_{2}(t)$ for the air-water interface.
+(a) The $C_{2}(t)$ for $d=1,4,6$ (Å) and the biexponential decay function $C_{2}(t)=A_{1}\exp(t/\tau_{1})+A_{2}\exp(t/\tau_{2})$.
+$d$-dependence of (b) $C_{2}(t^{*})$ for water molecules at the air-water interface with $t^{*}=1,2,5$ (ps); (c) relaxation time ($\tau_{1}$ and $\tau_{2}$), and (d) the amplitude ($A_{1}$ and $A_{2}$).
+Dotted line is a horizontal line for easier viewing.
+As $d$ increases to $d=4$ (Å), $C_{2}(t^{*})$ and $\tau_{i}$, $A_{i}$ all converge to a fixed value, respectively.
+(We also obtained consistent results for systems with 125 and 216 water molecules, please refer to Supporting Information.)
+As shown in Figure 5a, the orientation correlation function fits well with a bi-exponential decay function.
+The two relaxation rates ($1/\tau_{1}$ and $1/\tau_{2}$) differ; the larger relaxation time $\tau_{2}$ increases with the increase of $d$, while the smaller one, $\tau_{1}$, remains relatively unchanged with variations in $d$.
+In Figure 5b, the values of the orientational correlation function $C_{2}(t)=\langle P_{2}(\hat{u}(0)\hat{u}(t))\rangle$ at three reference time $t^{*}=1,2,5$ (ps) are plotted.
+At all these times, $C(t^{*})$ increases with increasing $d$ and no longer increases significantly for $d\geq 4$ (Å).
+Similar to the earlier results on HB population correlation functions and the HB reaction rates, as $d$ increases, $C_{2}(t^{*})$ and associated relaxation times converge to a fixed value, respectively, which characterizes the decay time of the orientation relaxation process of OH bonds.
+From the convergence trend of the relaxation times and corresponding amplitudes in Figure 5c,d, we find that at the interface with a thickness greater than 4 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water.
+Experimentally, the information on the interface molecules can be obtained from the SFG spectrum.
+The vibration relaxation time of the interface water molecules can be obtained based on the SFG spectrum using analysis techniques such as Singular Value Decomposition (SVD)ichi Inoue et al.
+(2020).
+We also defined the interface molecule orientation relaxation time from the $C_{2}(t)$ correlation of the interface water molecules.
+However, due to the arbitrariness of the specific form of defining the relaxation time, the results we obtained cannot be directly and accurately compared with the experimental results.
+Because of the current huge challenges in directly measuring water, we compared the data obtained from interfacial HB dynamics, interfacial OH orientation relaxation dynamics, and AIMD simulations.
+Our results from interfacial HB dynamics are consistent with the results from $C_{2}(t)$ and density profile from AIMD simulations of the air-water interface.
+Furthermore, the interface thickness aligns with values obtained through experimental measurements and AIMD simulations, as detailed in Table 1.
 Therefore, we arrive at a consistent conclusion on the issue of estimating the thickness of the air-water interface, from the perspective of HB dynamics and OH reorientation relaxation.
-Furthermore, the interface thickness aligns with values obtained through experimental measurements and AIMD simulations, as detailed in Table 1.
-Figure 5: The $d$-dependence of (a) $C_{2}(t^{*})$ and (b) $\tau_{2}(t^{*})$ for water molecules at the air-water interface with $t^{*}=1,2,5$ (ps), respectively.
-As $d$ increases to $d=4$ (Å), $C_{2}(t^{*})$ and $\tau_{2}$ converge to a fixed value, respectively.
-Table 1: The air-water interface thickness obtained by experiments and computer simulations ("-" denotes without using "10–90" thickness).
-Method Thickness Notation $T$ (K) $d$ (Å)
-Ellipsometry (Rayleigh) - 293.15 3.0
-Ellipsometry (Raman and Ramdas (1927)) - 293 5.0
-Ellipsometry (McBain, Bacon, and Bruce (1939)) - 293 $\geq$2.26
-Ellipsometry (Kinosita and Yokota (1965)) - 293 7.1
-X-Ray Reflectivity(Braslau et al. (1988)) - 298 3.24$\pm$ 0.05
-MD (TIP4P/FQ, Liu et alLiu, Harder, and Berne (2005)) 10–90 298.15 3.5
-MD (TIP4P-POL2, Kuo et alKuo et al. (2006)) 10–90 298 3.9
-MD (SPC-FW, Wick et alWick et al. (2007)) 10–90 298 3.2
-MD (D-C, Wick et alWick et al. (2007)) 10–90 298 3.2
-MD (TIP4P, Wick et alWick et al. (2007)) 10–90 298 3.4
-MD (SPC/E, Fan et alFan et al. (2009)) 10–90 300 6.0
-MD ( SPC/E, Willard and ChandlerWillard and Chandler (2010)) 10–90 298 8.0
-MD (SPC/E, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.27
-MD (TIP4P/2005, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.05
-CPMD/BLYP (Kuo et alKuo et al. (2006)) 10–90 300 1.7
-BOMD/BLYP-D (Kühne et al. Kühne et al. (2011)) 10–90 300 3.8
-BOMD/BLYP-D2 (Pezzotti et alPezzotti, Galimberti, and Gaigeot (2017) ) 10–90 315 3.5
-BOMD/BLYP-D3 (Jeon el al. Jeon et al. (2017)) 10–90 298 $\sim$3.0
-BOMD/BLYP-D3/LC&IHB (this work) - 300 4.0
+Table 1: Air-Water Interface Thickness Obtained by Experiments and Computer Simulations.
+Methods $T$ (K) $d$ (Å)
+ellipsometry (Rayleigh) 293.15 3.0
+ellipsometry (Raman and Ramdas (1927)) 293 5.0
+ellipsometry (McBain, Bacon, and Bruce (1939)) 293 $\geq$2.26
+ellipsometry (Kinosita and Yokota (1965)) 293 7.1
+X-ray reflectivity(Braslau et al. (1988)) 298 3.24$\pm$ 0.05
+DeePMD/MB-pol/Free OH (this work) 300 5.0
+BOMD/BLYP-D3/LC&IHB (this work) 300 4.0
+DeePMD/MB-pol/LC&IHB (this work) 300 4.0
+There is one more aspect worth noting when studying the air-water interface based on the aforementioned definition of H-bonds.
+Some complications arise in interpreting vibrational SFG spectroscopy results Tang et al.
+(2017).
+An alternative perspective focusing on nonbonded OH groups better establishes a direct correlation between free OH and the 3700 cm-1 peak, typically attributed to free OH groups Tang et al.
+(2017).
+Considering the dynamics of free OH groups, the thickness of the air-water interface is also estimated from simulations.
+Details on the methods and results can be found in Supporting Information.
+From the $d$-dependence of the interfacial free OH correlation function, we determined an interface thickness of approximately 5 Å.
+As shown in Table 1, this value slightly deviates from the main results presented in this paper, highlighting that the measured thickness of the air-water interface varies depending on the properties of interest.
+Thus, to accurately determine the water interface thickness, it is essential to integrate findings from these varied perspectives.
 IV Conclusions
-In this study, we have developed a "two-extremes" approach to investigate the HB dynamics at the air-water interface and to determine the interface’s thickness.
+In this study, we have developed a two-extremes approach to investigate the HB dynamics at the air-water interface and to determine the interface’s thickness.
 One extreme scenario underestimates the HB breaking rate constant, while the other overestimates it, implying that each scenario provides only a partial insight into the interfacial HB dynamics.
-Subsequently, based on the DFTMD simulations, we have applied our approach to two distinct system properties: HB relaxation and HB reaction rate constants at the air-water interface.
-Our results across all these properties indicate that the predictions from both extreme scenarios converge as the thickness of the air-water interface increases to 4.0 Å.
+Subsequently, based on DFTMD and DeePMD simulations, we have applied our approach to two distinct system properties: HB relaxation and HB reaction rate constants at the air-water interface.
+Our results across both properties indicate that the predictions from both extreme scenarios converge as the thickness of the air-water interface increases to 4 Å.
 Thus, we have reason to believe the thickness, which falls between these two extremes, converges at this critical value.
@@
 Beyond the scope of HB dynamics, this approach can be extended to other properties like molecular orientation distribution Rao, Tao, and Wang (2003); Gan et al.
-(2006b) and SFG spectrum Nihonyanagi et al.
+(2006b), free OH dynamics Tang et al.
+(2017), and SFG spectrum Nihonyanagi et al.
 (2011, 2015); Tang et al.
-(2020), as well as other solution interfacial surfaces where statistical properties of the interface and bulk phase differ significantly.
-Looking ahead, our study could also inspire researchers to examine the hydration shells of ions, and then study the effects of ions through HB dynamics.
+(2020), or other systems like solution interfacial surfaces where statistical properties of the interface and bulk phase differ significantly.
+Looking ahead, this study could inspire further research into ions’ hydration shells by examining ions’ effects through HB dynamics.
+It is important to note that our current approach focuses on probing the local environment by analyzing water pairs within HB dynamics, without considering the collective behaviors of water molecules in the HB networks.
+However, the collective dynamics of many water molecules have been increasingly discussed in recent research, demonstrating a close relationship with observable properties such as dielectric spectroscopy and time-dependent vibrational spectroscopy.Popov et al.
+(2016); Fecko et al.
+(2003, 2005); Hölzl, Forbert, and Marx (2021); Offei-Danso et al.
+(2023); Malosso et al.
+(2024) We hope our study will inspire further innovative ideas on related topics by incorporating considerations of the collective nature of water.
+Notes
+The authors declare no competing financial interest.
+Code availability
+The codes utilized in this study are publicly accessible on GitHub at https://github.com/hg08/hb_ihb.
 Acknowledgements.
@@
 2011113214).
-We thank Marialore Sulpizi for the useful discussion.
-The simulations were performed on the Mogon ZDV cluster in Mainz and on the Cray XE6 (Hermit) at the HRLS supercomputing center in Stuttgart.
-Appendix A Hydrogen bond correlation functions
-[…]
-We use a configuration $r(t)$ to denote the positions of all the atoms in the system at time $t$.
-Either of the criteria above allows one to define an HB population $h[r(t)]=h(t)$, which equals 1 when a tagged pair of molecules are H-bonded, and 0 otherwise.
-The fluctuation in $h(t)$ from its time-independent equilibrium average is defined by Chandler (1987) $\delta h=h(t)-\langle h\rangle$.
-The probability that a specific tagged pair of molecules is H-bonded in a large system is extremely small [^1] The average number of H-bonds in an equilibrium of $N$ water molecules is: $\frac{1}{2}N(N-1)\langle h\rangle$, i.e., $\frac{1}{2}N(N-1)\langle h\rangle\sim N$.
-It implies that $\frac{1}{2}N\langle h\rangle\sim 1$, or $\langle h\rangle\sim\frac{2}{N}$.
-Therefore, for a large system, $\langle h\rangle$ is extremely small., then $\delta h(t)=h(t)$.
-Therefore, the correlation of $\delta h(t)$ can be written as
-$\langle\delta h(0)\delta h(t)\rangle=\langle h(0)h(t)\rangle,$
-[…]
-The correlation function $c(t)=\langle h(0)h(t)\rangle/\langle h\rangle$ describes the structural relaxation of H-bonds Chandra (2000); Benjamin (2005).
-Here the average $\langle h\rangle$ of the HB population is the probability that a pair of randomly chosen water molecules in the system is H-bonded at any time $t$.
-The function $c(t)$ measures correlation in $h(t)$ independent of any possible bond-breaking events, and it relaxes to zero when $t$ is large Rapaport (1983).
-Because the thermal motion can cause distortions of H-bonds from the perfectly tetrahedral configuration, water molecules show a librational motion on a time scale of $\sim$ 0.1 ps superimposed to rotational and diffusional motions ($>1$ ps), which causes a time variation in interaction parameters.
-A new HB population $h^{\text{(d)}}(t)$ was also defined to obviate the distortion of real HB dynamics due to the above geometric definition Sciortino and Fornili (1989); Chandra (2000).
-It is 1 when the inter-oxygen distance of a particular tagged pair of water molecules is less than $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise.
-The H-bonds between a tagged molecular pair that satisfy the condition $h^{\text{(d)}}(t)=1$ may have been broken, but they may more easily form H-bonds again.
-The correlation function
-$n(t)=\langle h(0)[1-h(t)]h^{\text{(d)}}(t)\rangle/\langle h\rangle$ (12)
-represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules are unbonded but remain separated by less than $r_{\mathrm{OO}}^{\mathrm{c}}$ Chandra (2000).
-The rate of HB relaxation to equilibrium is characterized by the reactive flux Luzar (2000)
-$k(t)=-\frac{\mathrm{d}c(t)}{\mathrm{d}t},$ (13)
-which quantifies the rate that an initially present HB breaks at time $t$, independent of possible breaking and reforming events in the interval from 0 to $t$.
-Therefore, $k(t)$ measures the effective decay rate of an the initial set of H-bonds Chandler (1987).
-For bulk water, a $\sim 0.2$-ps transient period exists, during which $k(t)$ changes quickly from its initial value Starr, Nielsen, and Sta (2000).
-However, at longer times, $k(t)$ is independent of the HB definitions.
-The dynamic evolution of HB population correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, in the two scenarios, is shown in Fig.
-A6a and b, respectively.
-This figure applies the AHD criteria to H-bonds.
-A critical thickness $d_{f}$ (equal to or greater than 4.0 (Å)) exists.
-Beyond this thickness, the time evolution curves of the correlation function $c(t)$ and $c^{\rm(s)}(t)$ do not significantly change with further increases in $d$.
-A comparison between the two scenarios shows that in scenario 2, the correlation function $c^{\rm(s)}(t)$ decays faster than in scenario 1.
-This is visually indicated by a black directed line segment in the figure.
-This observation is consistent with conclusions drawn from another criterion, the ADH criterion, as we have shown in the main text.
-The result, where $c^{\rm(s)}(t)$ decays faster in scenario 2 compared to $c(t)$ in scenario 1, aligns with findings from the ADH standard.
-This suggests a level of agreement between the different criteria applied to analyze the air-water interface.
-Figure A6: The $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ obtained from the AHD criteria: (a) LC; (b) IHB. In scenario 1, the sampling is performed every 4 ps.
-Appendix B Hydrogen bond breaking and reforming rate constants
-Assume that each HB acts independently of other H-bonds Luzar and Chandler (1996b); Luzar (2000), and due to the detailed balance condition, one obtains $\tau_{\mathrm{HB}}={(1-\langle h\rangle)}/{k}$, where $k$ is the rate constant of breaking an HB, i.e., the forward rate constant Chandler (1986, 1978).
-Correspondingly, the backward rate constant $k^{\prime}$ is represented by the rate constant from the HB-on state to the HB-off state for a tagged pair of molecules.
-Based on the functions (see Appendix  A) $n(t)$, $h(t)$, $h^{\text{(d)}}(t)$, and $k(t)$, Khaliullin and Kühne (2013) have obtained the ratio $k/k^{\prime}$ of HB breaking and reforming rate constants in bulk water, and then the lifetime and relaxation time of H-bonds from simulations.
-Here, for the air-water interface, we obtain the optimal solution range of $k$ and $k^{\prime}$ from the relationship between the reactive flux $k(t)$ and the correlation functions $c(t)$ and $n(t)$:
-$k(t)=kc(t)-k^{\prime}n(t).$ (14)
-We obtain the optimal value of the rate constants, $k$ and $k^{\prime}$, by the least squares fit of $k(t)$, $c(t)$, and $n(t)$ beyond the transition phase.
-The function $c(t)$ is regarded as a column vector composed of $(c_{1},\cdots,c_{P})^{\mathrm{T}}$, and is denoted as ${\bf c}$, with $c_{i}$ representing the value of correlation $c(t)$ at $t=i$.
-Similarly, $n(t)$ and $k(t)$ can also be denoted as ${\bf n}$ and ${\bf k}$, respectively.
-Then, the rate constants $k$ and $k^{\prime}$ are determined from the matrix ${\bf A}=\begin{bmatrix}{\bf c}&{\bf n}\end{bmatrix}$:
-$\begin{bmatrix}k\\ -k^{\prime}\end{bmatrix}=({\bf A}^{\mathrm{T}}{\bf A})^{-1}{\bf A}^{\mathrm{T}}{\bf k}.$ (15)
-For bulk water and the air-water interface, the optimal $k$ and $k^{\prime}$ are reported in Table  B2 and  B3.
-Table B2: The rate constants $k$ and $k^{\prime}$ for the bulk water and the air-water interface (the time region $0.2<t<$ 2 (ps)).
-Criterion $k$ (b) [^2] The unit for $k$ ($k^{\prime}$) is ps-1.
-b: bulk; i: interface.
-$k^{\prime}$ (b) $\tau_{\mathrm{HB}}$ (b) [^3] The unit for $\tau_{\mathrm{HB}}=1/k$) is ps.
-$k$ (i) $k^{\prime}$ (i) $\tau_{\mathrm{HB}}$ (i)
-ADH 0.296 0.988 3.380 0.323 0.765 3.101
-AHD 0.288 1.149 3.470 0.314 0.887 3.184
-Table B3: The $k$ and $k^{\prime}$ for the bulk water and the air-water interface (the time region $122<t<$ 12 (ps)).
-Criterion $k$ (b) $k^{\prime}$ (b) $\tau_{\mathrm{HB}}$ (b) $k$ (i) $k^{\prime}$ (i) $\tau_{\mathrm{HB}}$ (i)
-ADH 0.115 0.039 8.718 0.157 0.068 6.372
-AHD 0.105 0.047 9.496 0.155 0.088 6.472
-Figure B7: The $d$-dependence of the rate constant $k^{\prime}_{\rm IHB}$ and $k^{\prime}_{\rm LC}$, obtained by the IHB (solid lines) and the Luzar-Chandler (dot-dashed lines) methods, respectively.
-The corresponding reaction rate $k^{\prime}_{\rm bulk}$ for bulk water is also represented as dashed lines.
-The ADH criterion of H-bonds is used and the fits are carried on the time region $120.2<t<$ 12 (ps).
-To obtain $k$ and $k^{\prime}$, we perform the fitting in short and long-time regions, respectively.
-Since HB dynamics show different relaxation behaviors in the time of $\sim 1$ ps and in that greater than $\sim 1$ psStarr, Nielsen, and Stanley (2000), we will fit the short and long time intervals respectively.
-We note that in the long time region ($122<t<$ 12 (ps)), the value of HB lifetime $\tau_{\mathrm{HB}}$ is larger than that in short one ($20.2<t<$ 2 (ps)), no matter for the bulk water or the air-water interface.
-A larger $\tau_{\mathrm{HB}}$ value means that the distance between a pair of water molecules stays within $r_{\mathrm{OO}}^{c}$ for a longer time.
-Appendix C Details of Scenario 1: the LC method
-In the LC scenario, to obtain interface HB dynamics, we use molecule sampling at the instantaneous interface.
-Let $T_{\text{e}}$ be the time it takes for all water molecules in the simulation box to traverse the interface and bulk phase.
-If the trajectory length $t_{\text{traj}}$ satisfies the condition $\tau_{\text{HB}}\ll t_{\text{traj}}\ll T_{\text{e}}$, we sample time points $t_{i}$ $(i=1,\cdots,M)$ that are evenly spaced on the trajectory, then obtain interface HB dynamics using the following procedure:
-  1. 1.
-    For each sampling time $t_{i}$, define an interface with a thickness $d$ (see Fig.
-1).
-Then, select a set $S_{i}$ of water molecules in the interfacial layer $\mathcal{I}(d;t)$, i.e., $\mathbf{r}(t_{i})\in\mathcal{I}(d;t)$, and calculate the correlation functions $c(t)$, $n(t)$ and $k(t)$ through $t_{\text{traj}}$ for these water molecules.
-  2. 2.
-    Calculate average functions of the correlation functions $c(t)$, $n(t)$, and $k(t)$ respectively over all sampling time points and all trajectories.
-  3. 3.
-    Calculate reaction rate constants of breaking and reforming in the interfacial layer $\mathcal{I}(d;t)$ by Eq. 14.
-In scenario 1, the sampling is performed every 4 ps.
-Appendix D Rotational anisotropy decay of OH stretch at the water/vapor interface
-We examine the results for the water interface thickness from the interfacial HB dynamics by polarization anisotropy decay of OH stretch.
-The polarization anisotropy decay of the OH stretch can also provide information on the dynamics of water  Woutersen and Bakker (1999); Cowan et al.
-(2005); Kraemer et al.
-(2008); Yagasaki, Ono, and Saito (2009).
-We assume that the anisotropy decay $R(t)$ is only due to the orientational relaxation of water molecules.
-Within the Condon approximation  Schmidt, Corcelli, and Skinner (2005), it is directly related to the orientational correlation function $C_{2}(t)$ Tokmakoff (1996); Rezus and Bakker (2006); Yagasaki, Ono, and Saito (2009); Bakker and Skinner (2010) through $R(t)=\frac{2}{5}C_{2}(t)$.
-The $C_{2}(t)$ is given by the rotational time-correlation function
-$C_{2}(t)=\langle P_{2}(\hat{u}(0)\cdot\hat{u}(t))\rangle,$ (16)
-[…]
-For the air-water interface, using the method in section II, we first obtain the instantaneous interface with thickness $d=1,\cdots,6$ (Å).
-Then, for each value of $d$, the LC method is used to obtain $C_{2}(t)$ for water molecules at the air-water interface.
-Here we introduce a fitting with an exponential and provide a decay time $\tau_{2}$, i.e., $C_{2}(t)=Ae^{-t/\tau_{2}}$, with $A=1.0$ for different thickness values, and the $d$-dependence of $C_{2}(t^{*})$ and $\tau_{2}(t^{*})$ is shown in Fig.
-5.
-When the interface thickness $d$ is smaller than 4.0 (Å), the decay time $\tau_{2}$ increases linearly with $d$.
-The OH orientational relaxation time $\tau_{2}$ for the neat water surface is from $1.4$ to $2.0$ (ps), with std.
-error $\delta d=0.1$ (ps), which is consistent with the vibrational relaxation time of the free OH for the neat water ichi Inoue et al.
-(2020).
-Similar to the correlation functions, $c(t)$ and $k(t)$, for interfacial water molecules, as $d$ increases, $\tau_{2}$ converges to a fixed value, which characterizes the decay time of the orientation relaxation process of OH bonds in bulk water.
-From the convergence trend of $\tau_{2}$ in Fig.
-5(b), we find that at the interface with a thickness greater than 4.0 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water.
-Appendix E Computational Methods
-To describe the subtleties of H-bonding in water Laasonen et al.
-(1993), we have performed DFTMD simulations Marx and Hutter (2000) for bulk water and the air-water interface.
-The simulations make use of technologies that have been successfully tested on water and solutions Khatib et al.
-(2016a, b); Khatib and Sulpizi (2017), namely the Goedecker-Teter-Hutter (GTH) pseudopotentials Goedecker, Teter, and Hutter (1996); Hartwigsen, Goedecker, and Hutter (1998); Krack (2005), Generalised Gradient Approximation (GGA) of the exchange-correlation functional Becke (1988); Lee, Yang, and Parr (1988), and dispersion force correction, DFT-D3 Grimme et al.
-(2010); Klimeš and Michaelides (2012).
-By eliminating the strongly bound core electrons, the GTH pseudopotentials reduce the number of occupied electronic orbitals that have to be treated in an electronic structure calculation.
-There are dual-space Gaussian-type pseudopotentials that are separable and satisfy a quadratic scaling to system size Lu et al.
-(2019).
-The GGA functionals generally describe the dipole and quadruple moments of the molecules quite well, and the DFT-D3 correction treats the van der Waals dispersion forces and improves the structural properties without more computational cost and thus can be used for the air-water interface.
-Figure E1: The partial radial distribution functions (RDFs) for the simulated bulk water system.
-Figure E2: Probability distribution $P(z)$, along the $z$-axis, of H2O in the slab.
-[…]
-The bulk water system consisted of 128 water molecules in a periodic box of size $15.64\times 15.64\times 15.64$3, and with a density of 1.00 g cm-3.
-The slab consisted of 128 water molecules in a periodic box of size $15.64\times 15.64\times 31.28$3.
-The length of each trajectory is $60$ (ps).
-The RDFs $g_{\mathrm{OO}}(r)$ and $g_{\mathrm{OH}}(r)$ for the bulk water systems are shown in Fig.
-E1.
-The probability distribution of O and H atoms in the simulated model of the air-water interface is shown in Fig.
-E2.
+G.H.
+thanks Yuliang Jin, Rémi Khatib and Marialore Sulpizi for the useful discussions.
+The DFTMD simulations were conducted on the Mogon ZDV cluster in Mainz and the Cray XE6 (Hermit) at the HRLS supercomputing center in Stuttgart.
+The DeePMD simulations were performed on Triton, which is provided by the Aalto Science-IT project.
+Supporting Information
+The Supporting Information is available free of charge at https://pubs.acs.org/doi/10.1021/acs.jctc.4c00457.
+Hydrogen bond correlation functions, HB breaking and reforming rate constants, details of Scenario 1: the LC method, rotational anisotropy decay of OH stretch at the air-water interface, computational methods including DFTMD and DeePMD simulations, results based on DFTMD simulations for systems in other sizes, results based on DeePMD simulations for systems in other sizes and perspectives derived from the correlation function of free OH groups (PDF)